OpenStudy (anonymous):

the formula for the area of a triangle is a =1/2bh where b is the base of the triangle and h is the height of the triangle .what is the length of the base if the area is 22 cm to the 2 power and the height is 8 cm A. 4.5 B. 5 C. 5.5 D. 6

OpenStudy (anonymous):

CAN YYOU PLEASE EXPLAIN HOW TO DO IT

OpenStudy (anonymous):

ok. the area is given as \[22 cm ^{2}\] = (1/2) * base * height

OpenStudy (anonymous):

22 = (1/2)*base*8 Therefore, base = 2*22 / 8 = 44/8 = 22/4 = 5.5 cm

OpenStudy (anonymous):

ok thank you